OpenStudy (anonymous):

What is the solution set for |3x + 12|+ 5 = 23?

OpenStudy (anonymous):

@Lizzyeb631 first subtract 5 from both the sides tell us what will u get?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

@Lizzyeb631 u will get \[\left| 3x+12 \right| = 18\] pls do write the full equation from now on

OpenStudy (anonymous):

i got X=2 but how do i get a second answer do i plug it back in

OpenStudy (anonymous):

look at the equation there is a modulus when u remove the modulus sign from one side , "±" comes on the other side for example if \[\left| 7x + 16 \right| = 30\] then 7x + 16 = ± 30

OpenStudy (anonymous):

@Lizzyeb631 did u get it?

OpenStudy (anonymous):

yeah i got it know thx for help

OpenStudy (anonymous):

:)